OpenStudy (joshoyen):

Solve each equation. Check each solution. 1. - 4/x+1 = 5/3x+1

OpenStudy (joshoyen):

\[-\frac{ 4 }{ x+1 }=\frac{ 5 }{ 3x+1 }\]